# LoRE-Merging: Exploring Low-Rank Estimation For Large Language Model Merging

### v1

@@ 1 Introduction
To this end, we first empirically validate that task vectors exhibit a small number of dominant singular values, with the remaining singular values being significantly smaller in magnitude, as shown in Figure 1. Additionally, the dimension of the intersection of the images of two matrices is bounded by the minimum of their ranks. Therefore, we propose LoRE-Merging, a unified framework for model merging based on Low-Rank Estimation of task vectors, which eliminates the need for access to the base model. Specifically, given a set of FT models, we formulate the merging problem as an optimization problem whose goal is to simultaneously identify an approximate base model integrated with a set of low-rank task vectors that collectively approximate the behavior of the FT models. By leveraging low-rank estimations, task vectors become inherently less susceptible to interference, effectively addressing a fundamental challenge in model merging. We conduct extensive experiments on optimization modeling problems and math word problems to confirm the effectiveness of our method.
@@ 2 Related Works
Merging fine-tuned models has been shown to offer several benefits, such as improving performance on a single target task (Gupta et al., 2020; Choshen et al., 2022; Wortsman et al., 2022), enhancing out-of-domain generalization (Cha et al., 2021; Arpit et al., 2022; Ilharco et al., 2022; Ramé et al., 2023), creating multi-task models from different tasks (Jin et al., 2022; Li et al., 2022; Yadav et al., 2024), supporting continual learning (Yadav and Bansal, 2022; Yadav et al., 2023), and addressing other challenges (Don-Yehiya et al., 2022; Li et al., 2022). Among these methods, task-vector-based merging approaches play an important role. Task Arithmetic (Ilharco et al., 2022) first introduced the concept of task vectors and shows that simple arithmetic operations can be performed to obtain the merged models. Building on this idea, methods like DARE (Yu et al., 2024b) and Ties (Yadav et al., 2024) adopt pruning-then-scaling techniques to merge task vectors, based on the assumption that not all parameters equally contribute to the final performance. However, these methods based on sparsity estimation consistently suffer from the interference among task vectors and require access to the base model, thus limiting their overall effectiveness.
@@ 3.2 Implicit Low-Rank Estimation for Model Merging
As demonstrated in Yu et al. (2024b, a), model pairs would exhibit limited mergeability, especially when comprehensive fine-tuning or extended pre-training procedures are employed and result in substantial parameter shifts. Under such conditions, existing task-vector based merging methods struggle to work due to the significant representational divergence between the base model and its fine-tuned derivative. To address this challenge, we propose an implicit low-rank estimation model merging method, named LoRE-Merging, which not only employs the robustness of low-rank estimation against the interference but also eliminates the need for access to the base model.
@@
4 Experiments
Method Deepseek & NuminaMath WizardLM & WizardMath Checkpoints Avg.
GSM8K MATH GSM8K MATH EasyLP ComplexLP NL4OPT
Baseline 76.3 55.8 54.8 12.4 81.9 39.3 94.0 59.21
Average 75.0 45.8 58.8 12.6 75.9 40.3 91.6 57.14
DARE 81.0 54.2 14.9 3.7 80.7 35.1 95.1 52.10
Ties-Merging 80.8 51.6 58.5 11.8 82.4 42.7 94.8 60.37
[…]
Table 1: The evaluation results on math word problems and optimization modeling problems. We use the best performance of base models as the baseline.
Baselines & Settings
Evaluation
We first evaluate LoRE-Merging on math word problems using the popular benchmarks GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., ). For comprehensive evaluation, we test both DeepSeek-series models (NuminaMath-7B (Beeching et al., 2024) and DeepSeek-Math-7B-Base (Shao et al., 2024)) and LLaMA-series models (WizardLM-13B (Xu et al., 2023) and WizardMath-13B (Luo et al., 2023)). Additionally, we also evaluate the effectiveness of LoRE-Merging on another advanced task, i.e. mathematical optimization modeling problems (Ramamonjison et al., 2023; Huang et al., 2024, 2025). This task aims to generate solvable mathematical models given an optimization problem in natural language. As the lack of public models on this task, we first fine-tune Qwen-2.5-Coder-7B-Instruct model (Hui et al., 2024) with the datasets provided by Huang et al. (2025) and merge the checkpoints in the training process. The evaluations are conducted on MAMO dataset (Huang et al., 2024) which includes two subsets EasyLP and ComplexLP, and NL4OPT dataset (Ramamonjison et al., 2023).
Main Results
As shown in Table 1, LoRE-Merging achieves superior performance across most metrics, as well as the highest overall score. For the math word problem evaluation, our method demonstrates consistently superior performance compared to the baselines, except for the evaluation on the MATH dataset when merging DeepSeek-Math with NuminaMath. due to the significant performance gap between the base models, where DeepSeek-Math achieves only a score of 36.2 on the MATH dataset, while NuminaMath reaches 55.8. As indicated in Yao et al. (2024), a large performance gap can significantly impact the effectiveness of model merging. Another worthy-noting observation is that DARE demonstrates significantly poorer performance when merging WizardLM and WizardMath. This can likely be attributed to the substantial parameter divergence between these models, which results in the failure of calculating the task vector derived from the base model. In contrast, our LoRE-Merging with the approximate base model and low-rank task vectors demonstrates superior robustness and effectiveness in solving math word problems. For the evaluations on optimization modeling with checkpoints merging, we can see existing task-vector based merging methods consistently improve the performance because of the marginal gap between the checkpoints. Therefore, we believe that checkpoint merging can serve as a highly effective technique complementary to training methods, particularly our LoRE-Merging method. We also conduct a detailed analysis how our method enhance the modeling capacity on ComplexLP dataset. We found that the earlier checkpoint is more good at identifying the variables and parameters in the questions while the later one focuses on more complex components, such as formulating variables and the constraints. With the merging of task vectors, the merged model exhibits superior overall performance on the task.
5 Conclusion
In this paper, we propose a unified framework for merging homogeneous models based on low-rank estimation, named LoRE-Merging. The main motivation of our work is to estimate task vectors using low-rank approximation without the need of access to the base model. We achieve it by formulating the merging problem as an optimization problem. Extensive experiments demonstrate the efficacy and efficiency of our proposed methods.
Limitations

### v2

@@ 1 Introduction
To this end, we first empirically validate that task vectors exhibit a small number of dominant singular values, with the remaining singular values being significantly smaller in magnitude, as shown in Figure 1. Additionally, the dimension of the intersection of the images of two matrices is bounded by the minimum of their ranks. Therefore, we propose LoRE-Merging, a unified framework for model merging based on Low-Rank Estimation of task vectors, which eliminates the need for access to the base model. Specifically, given a set of FT models, we formulate the merging problem as an optimization problem whose goal is to simultaneously identify an approximate base model integrated with a set of low-rank task vectors. Together, these vectors collectively approximate the behavior of the FT models. By leveraging low-rank estimations, task vectors become inherently less susceptible to interference, effectively addressing a fundamental challenge in model merging. We conduct extensive experiments on optimization modeling problems and math word problems to confirm the effectiveness of our method.
@@ 2 Related Works
Merging fine-tuned models has been shown to offer several benefits, such as improving performance on a single target task (Gupta et al., 2020; Choshen et al., 2022; Wortsman et al., 2022), enhancing out-of-domain generalization (Cha et al., 2021; Arpit et al., 2022; Ilharco et al., 2022; Ramé et al., 2023), creating multi-task models from different tasks (Jin et al., 2022; Li et al., 2022; Yadav et al., 2024), supporting continual learning (Yadav and Bansal, 2022; Yadav et al., 2023), and addressing other challenges (Don-Yehiya et al., 2022; Li et al., 2022). Among these methods, task-vector-based merging approaches play an important role. Task Arithmetic (Ilharco et al., 2022) first introduced the concept of task vectors and shows that simple arithmetic operations can be performed to obtain the merged models. Building on this idea, methods like DARE (Yu et al., 2024b) and TIES-Merging (Yadav et al., 2024) adopt pruning-then-scaling techniques to merge task vectors, based on the assumption that not all parameters equally contribute to the final performance. However, these methods based on sparsity estimation consistently suffer from the interference among task vectors and require access to the base model, thus limiting their overall effectiveness.
@@ 3.2 Implicit Low-Rank Estimation for Model Merging
In this study, drawing upon methodologies similar to those presented by Matena and Raffel (2022), we investigate the model merging problem without presupposing specific characteristics of, or requiring access to, a base model. This methodological decision is underpinned by several key rationales. Firstly, in the context of checkpoint merging (Liu et al., 2024), a prevalent scenario involves access restricted solely to checkpoints saved during the training trajectory, before the finalization of a base model. Consequently, in such instances, the assumption of a pre-defined base model is untenable. Furthermore, as demonstrated by Yu et al. (2024b, a), model pairs frequently exhibit limited mergeability, particularly when subjected to extensive fine-tuning or prolonged pre-training, which can induce substantial parametric shifts. Under these circumstances, existing task-vector-based merging techniques often prove less effective due to significant representational divergence between an original base model and its fine-tuned counterpart. To surmount this challenge, we introduce LoRE-Merging, an implicit low-rank estimation approach to model merging. This method leverages the inherent robustness of low-rank estimation against perturbations while obviating the requirement for base model access.
@@
4 Experiments
Method DPSK & Numina LM & Math Math & Code Checkpoints Merging Avg.
GSM8K MATH GSM8K MATH MMLU GLUE MBPP EasyLP ComplexLP NL4OPT
Baseline 76.3 55.8 54.8 12.4 52.0 63.3 32.0 81.9 39.3 94.0 56.18
Average 75.0 45.8 58.8 12.6 52.8 61.7 28.0 75.9 40.3 91.6 54.25
DARE 81.0 54.2 14.9 3.7 52.7 59.1 27.6 80.7 35.1 95.1 50.41
TIES 80.8 51.6 58.5 11.8 53.1 59.3 26.8 82.4 42.7 94.8 56.18
[…]
Table 1: Evaluations on various benchmarks. LM and Math are Wizard-series models, namely WizardLM-13B and WizardMath-13B. Code is llama-2-13b-code-alpaca model. The score of baseline is the higher one of base models.
Datasets $\mu=0$ $\mu=0.01$ $\mu=0.1$ $\mu=1.0$
GSM8K (%) 81.3 82.0 79.9 67.3
MATH (%) 53.8 54.5 53.8 42.4
Table 2: The ablation study for the hyperparameter $\mu$ (with $\lambda=1.0$) on DPSK & Numina.
Datasets $\lambda=0.5$ $\lambda=1.0$ $\lambda=1.5$
GSM8K (%) 18.9 82.0 79.1
MATH (%) 33.1 54.5 51.0
Table 3: The ablation study for the hyperparameter $\lambda$ (with $\mu=0.01$) on DPSK & Numina.
Baselines & Settings
Evaluation
We first evaluate LoRE-Merging on diverse benchmarks, including GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., ) (math word problem), MMLU (Hendrycks et al., ), GLUE(Wang et al., 2019) (commonsense reasoning) and MBPP(Austin et al., 2021) (code task). We evaluate DeepSeek-series models (NuminaMath-7B (Beeching et al., 2024) and DeepSeek-Math-7B-Base (Shao et al., 2024)) and LLaMA-series models (WizardLM-13B (Xu et al., 2023), WizardMath-13B (Luo et al., 2023) and LLaMA-2-13B-Code model). Additionally, we also evaluate on the advanced task, i.e. mathematical optimization modeling problems (Ramamonjison et al., 2023; Huang et al., 2024, 2025). This task aims to generate solvable mathematical models given an optimization problem in natural language. As the lack of public models on this task, we first fine-tuned Qwen-2.5-Coder-7B-Instruct model (Hui et al., 2024) with the dataset provided by Huang et al. (2025) and merge checkpoints in the training process. The evaluations are conducted on MAMO dataset (Huang et al., 2024) which includes two subsets EasyLP and ComplexLP, and NL4OPT dataset (Ramamonjison et al., 2023).
Main Results
As shown in Table 1, LoRE-Merging achieves superior performance across most metrics, as well as the highest overall score. For the math word problem evaluation, our method demonstrates consistently superior performance against baselines, except for the evaluations on MATH (DPSK & Numina) and MBPP datasets. We think this is because of the significant performance gap between the base models, where DeepSeek-Math achieves only a score of 36.2 on the MATH dataset, while NuminaMath reaches 55.8. As indicated in Yao et al. (2024), a large performance gap can significantly impact the effectiveness of model merging. Another worthy-noting observation is that DARE demonstrates significantly poorer performance when merging WizardLM and WizardMath. This can likely be attributed to the substantial parameter divergence between these models, which results in the failure of calculating the task vector derived from the base model. In contrast, our LoRE-Merging with the approximate base model and low-rank task vectors demonstrates superior robustness and effectiveness in solving math word problems. For the evaluations on optimization modeling with checkpoints merging, we can see existing task-vector based merging methods consistently improve the performance because of the marginal gap between the checkpoints. Therefore, we believe that checkpoint merging can serve as a highly effective technique complementary to training methods, particularly our LoRE-Merging method. We also conduct a detailed analysis how our method enhance the modeling capacity on ComplexLP dataset. We found that the earlier checkpoint is more good at identifying the variables and parameters in the questions while the later one focuses on more complex components, such as formulating variables and the constraints. With the merging of task vectors, the merged model exhibits superior overall performance on the task.
Ablations
We conduct a systematic empirical analysis of the selection of hyperparameters $\lambda$ and $\mu$, as presented in Table 3 and Table 3. Our results show that the best performance is achieved with $\lambda=1.0$ and $\mu=0.01$. Notably, variations in the hyperparameters around these values do not significantly impact the final performance, indicating the robustness of LoRE-Merging.
5 Conclusion
In this paper, we propose a unified framework for merging models based on low-rank estimation, named LoRE-Merging. We achieve it by formulating the merging problem as an optimization problem. Extensive experiments demonstrate the efficacy and efficiency of our proposed methods.
Limitations
Appendix A Appendix
Method Average TIES-Merging Twin-Merging LoRE-Merging
Acc. on GSM8K 75.0 80.8 79.9 81.0
Runtime 4.2s 5min 29s 17min 44s 12min 24s
Table 4: Time cost comparisons among different merging methods.
A.1 Speed and Computational Cost
While standard SVD exhibits computational inefficiency for extremely large matrices comprising billions of elements, its application to LLM presents a substantially different computational profile. Despite LLMs containing billions of parameters in aggregate, SVD operations are performed on individual parameter matrices, each typically comprising only millions of entries. For instance, in the Qwen2.5-72B architecture, the largest matrix requiring decomposition is dimensioned at $8192\times 28564$, while for Qwen2.5-7B, the corresponding matrix has dimensions of $3854\times 18944$. Thus, the substantial parameter differential between LLM scales does not translate to proportionally expanded matrix dimensions. In our implementation, merging operations for 7B-scale models require approximately 5 minutes using Ties-Merging, while LoRE-Merging necessitates approximately 12 minutes. However, compared to another SVD-based mering method, like Twin-Merging (Lu et al., 2024), our method exhibit superior performance on efficiency.
A.2 Task Vector Rank Validation
In this subsection, we validate the low-rank properties underlying the low-rank assumption. Specifically, we focus on the checkpoint merging problem and compute the rank of the task vectors. As previously discussed, we set the rank $r$ as $r=0.2\times\min\{m,n\}$ for any given task vector $\bm{\delta}$.
The distribution of the largest 100 singular values for Layer $1$ is presented in Figure 1. Our experimental results reveal that $\sigma_{r}\leq 0.05\times\sigma_{1}$, indicating that the singular values set to 00 in low-rank estimation are significantly smaller than the largest singular value across all linear layers. This finding supports the validity of adopting a low-rank approximation for task vectors, as it reflects the inherent structure of the data.